OpenStudy (anonymous):

Andrew purchased a living room set for $2,276 using an 18-month deferred payment plan with an interest rate of 23.49%. What is the balance after the deferment period if no payments are made? $2,276.00 $2,810.63 $3,123.34 $3,226.39

OpenStudy (anonymous):

That's a high interest rate. Okay here we go: Financial Calculator PV = $2,276 n = 1.5 (18 months converted to years) i = 23.49 which gives you a future value of $3,123.34 (rounded to two decimal places) Formula: \[FV = P(1 + \frac{ r }{n })^{nm} \] Where: FV= future value p = Principle (present value) r = interest rate n = compounding periods per year m = number of years